OpenStudy (anonymous):

expression equal to 3√ 48 48 2 3√6 6 3√2 12 3√2

OpenStudy (zzr0ck3r):

\(3\sqrt{48}=3\sqrt{16*3}=3\sqrt{16}\sqrt{3}= \ ?\)

OpenStudy (cookie_2046):